solve by factoring
2y^2+y-6=0

Answer by unlockmath(1688)

Hello,
This we can be factored as:
(2x-3)(x+2)=0
Solve for x to get:
2x-3=0
Add 3 and divide by 2 to get:
x=2/3
The other is:
x=-2
(Hint: factor means 2 things being multiplied together.
Make sense?
RJ
www.math-unlock.com